The units digit of a two digit number is 1 more than twice the tens digit. If the digits are reversed,the new number is 36 more than the original number. Find the number.
Answer by rimjhim.tomar(21)

let the tens digit be x
hence units digit would be 2x+1
and the two digit number would be =10(tens digit)+unit digit
=10(x)+2x+1
now,
when the digits are reversed,
we have-
the tens digit of the new number=(unit digit of the original number)= 2x+1
and similarly,unit digit of the new number=x
hence the new number=10(2x+1)+x
given that-
the new number is 36 more than the original number
or
10(x)+2x+1+36=10(2x+1)+x
10x+2x+1+36=20x+10+x
12x+37=21x+10
12x-21x=10-37
-9x=-27
x=(-27)/(-9)
x=3
the number= 10(x)+2x+1
=10(3)+2(3)+1
=30+6+1
=37